OpenStudy (anonymous):

find the volume. a cone with a radius of 20mm and a height of 24mm

OpenStudy (hockeychick23):

V= πr^2(h/3) V=π(20^2)(24/3) V=10053.1

OpenStudy (anonymous):

V= πr^2(h/3) V=π(20^2)(24/3) V=10053.1 10053.1mm

OpenStudy (anonymous):

how do you find out if was h/3

OpenStudy (hockeychick23):

h/3 is always a part of the equation to find the volume of a cone. The volume of a cone is found by using the formula πr^2(h/3)

OpenStudy (anonymous):

That is correct what @hockeychick23 said. It's part of the formula.

OpenStudy (anonymous):

so you would just plung in the missing numbers and work it on from there

OpenStudy (anonymous):

Yes

OpenStudy (hockeychick23):

correct :)

OpenStudy (anonymous):

oh ok so if i have diameter of 22 and a height of 16ft. it would b the same as last time?

OpenStudy (hockeychick23):

well for that one you need to divide the diameter by 2 to get the radius and use the same formula. so your radius is 11 for this one.

OpenStudy (anonymous):

Radius is Diameter/2. So the radius would be 11.

OpenStudy (anonymous):

so you would multiply the height by 3.14

OpenStudy (hockeychick23):

correct :) so you do: V= π11^2(16/3) V=π 121(16/3) V=2027.37

OpenStudy (anonymous):

what if they only give u a radius ?

OpenStudy (hockeychick23):

you cant solve it without a radius and a height if you are given a cone unless they give you the volume

OpenStudy (hockeychick23):

or the diameter instead of the radius, that works just as well

OpenStudy (anonymous):

it say a sphere with a radius of 16cm?

OpenStudy (hockeychick23):

sphere volume is a different equation, for a sphere you use: V= 4/3πr^3

OpenStudy (anonymous):

opps i mean 18

OpenStudy (hockeychick23):

Ok no problem, so V=4/3π(18^3)

OpenStudy (hockeychick23):

you will then get 24,429.02

OpenStudy (anonymous):

oh ok so it would b v=4/3 = 1*3.14(18^3)

OpenStudy (hockeychick23):

correct!

OpenStudy (anonymous):

oh ok. so just plug in the numbers then multiple it by 3.14?

OpenStudy (hockeychick23):

yepp :)

OpenStudy (anonymous):

oh ok i think i got it now thank you!!!

OpenStudy (hockeychick23):

sure!(: